The City of Denver, Colorado Recently pulled 525 city residents (R) are and non resident (R') about its proposal to install Smart Kiosks and increase the Parking Rates in the city during business hours. For each person polled, they recorded whether he or she was in favor of (F) or opposed to (F') the proposal.

Opinion
Residence In Favor (F) Opposed (F') Total
Residence (R) 90 120 210
Non Residence (R') 135 180 315
Total 225 300 525

a. What is the probability that a person selected at random is in favor of the proposal?

b. What is the probability that a person selected at random is in favor of the proposal, Given that he or she is a resident of Denver?

c. Determine whether the events "in favor" and "resident" are independent. Explain.
Answer #1

a)  probability that a person selected at random is in favor of the proposal = 225/525 = 0.4286

b) probability that a person selected at random is in favor of the proposal, Given that he or she is a resident of Denver

=90/210 = 0.4286

c) P(F|R) = 0.4286

P(F) = 0.4286

since, P(F|R) = P(F)

so, the events "in favor" and "resident" are independent